OpenStudy (anonymous):

Use the substitution method to solve the following system of linear equations: x + y = 6 x = –y + 2

OpenStudy (anonymous):

we have given that x + y = 6 and x = -y+2 substitute x = -y+2 in the first equation. so we get -y + 2 +y = 6 2 = 6 and this is not possible. So the system of given equations is inconsistent.